# Uranus and Neptune as methane planets: producing icy giants from refractory planetesimals

### v1

@@ 1 Introduction
The first models of Uranus and Neptune were constructed based on the premise that they were formed in the outer parts of a primitive proto-planetary nebula, where the elemental abundances were considered solar. Thus, assuming chemical equilibrium, early studies used calculations of the expected relative molecular abundances to gauge the potential composition of the planets that formed in this region (Reynolds and Summers, 1965; Podolak and Cameron, 1974). Ices, such as HO, CH, and NH were collectively found to be about twice as abundant as rocky materials. Hereafter, we refer to rocks as those refractory minerals in which the ratio of C to silicate-forming elements such as Si, Fe, Mg or Ca is less than 0.1, whereas in refractory organics (CHONs) this ratio is larger than 10 (Fomenkova et al., 1992). These early models typically considered a layered structure, with a rock core, an ice mantle, and an envelope rich in hydrogen and helium (Podolak and Reynolds, 1984; Podolak et al., 1995). Rocks were modeled mainly as silica (SiO), but sometimes mixed with other molecules (e.g., MgO, Fe, Ni) and water was considered the most abundant and therefore primary ice. The gas envelope was modeled mainly as a H+He solar mix with an admixture of HO, CH, and NH. Later developments highlighted the role of a more complex internal structure including material mixing and gradients (Helled et al., 2020; Helled and Fortney, 2020). There is increasing evidence however that the constituent building blocks of the outer planets might be refractory-rich, rather than water-rich, challenging common convention with respect to Uranus and Neptune. We summarize these arguments below, indicating that refractory materials in the outer solar system could be 3-6 times more abundant than water.
@@
Large KBOs like Pluto are of such considerable size, that they contain very little residual porosity. Thus, its measured bulk density of 1854 kg m (Stern
et al., 2015) directly translates into a similarly rock-rich composition. Other large KBOs like Eris and Haumea often have even higher densities, which imply yet richer rocky compositions (although these are typically associated with further mantle-striping modifications via giant collisions (Lissauer and
Stewart, 1993; Leinhardt et al., 2010; Barr and
Schwamb, 2016)).
Comets are nowadays regarded as highly refractory-rich bodies. There is a growing body of evidence that strongly suggest comets have refractory to ice mass ratios of at least 3, and potentially even twice that value (Rotundi et al., 2015; Fulle et al., 2016, 2017; Fulle
et al., 2019; Choukroun
et al., 2020).
Polluted white dwarfs are astrophysical mass spectrometers. Their atmospheres probe the composition of accreted exo-Solar planetary debris, inferring the bulk chemical composition of the pollutants (Hollands et al., 2018; Harrison et al., 2018, 2021). The typical finding is of almost purely rocky debris (Jura and
Xu, 2012; Jura and
Young, 2014; Pietro Gentile Fusillo
et al., 2017). It must be born in mind however that planetesimals can originate from different regions of the disk, devolatilize before their host star becomes a white dwarf (Malamud and
Perets, 2016, 2017a, 2017b), and the progression of accretion can complicate the association between pollution and pollutant (Brouwers
et al., 2022). Despite these complications, relatively rare few individual cases point to polluting debris which probably originate in the outer parts of their systems and which contain a significant mass fraction of water (Farihi et al., 2013; Raddi
et al., 2015; Xu et al., 2017). Yet even in these outlier cases the water fraction is 26%-38%, showing that a large rock/water mass ratio might be a ubiquitous feature of outer planetary disks.
@@ 2 Expected composition of outer solar system planetesimals
Present-day comets typically contain tiny amounts of CO and other hypervolatiles, relatively little water, and are primarily composed of refractory materials (Bockelée-Morvan and Biver, 2017; Fulle et al., 2017; Fulle
et al., 2019; Choukroun
@@
et al., 2020). Observations of various proto-planetary discs indirectly imply that when sufficiently far out in the disc, carbon is depleted from the gas phase and instead enters into solid grains locked inside planetesimals (McClure, 2019; Tabone et al., 2023). Below we however discuss direct observations of cometary dust from our own solar system, pointing towards a composition rich in carbon-bearing organic refractories. We show that between 1/3-1/2 of the mass of their refractory material is contained in organic rather than silicate material:
@@
Comet P/Halley was explored by Giotto, a flyby mission in 1986 (Keller et al., 1986), showing that cometary dust is extremely rich in organics relative to typical chondrites, and more similar to dust in dense molecular clouds. High-velocity collisions of dust particles with metal targets on the spacecraft provided a source of ions for mass-spectrometric analysis (Kissel
et al., 1986; Lawler and
Brownlee, 1992). Particles containing higher fractions of organics were relatively more abundant closer to the nucleus (Fomenkova, 1999), since organic components are more easily dissociated in the coma (Fomenkova et al., 1994; Wooden et al., 2007), providing an extended source for various gaseous species (Cottin and
Fray, 2008). Studies also discuss the detailed distribution of carbon among different compounds (Fomenkova, 1999). Overall, the mass ratio of rocks to organics in dust is between 2 and 1 (Fomenkova and
Chang, 1993). I.e, carbon-rich organics constitute as 33-50% of the dust mass.
@@
Comet 67P/Churyumov-Gerasimenko was explored by ESA’s Rosetta mission, which was the first to successfully rendezvous with a comet nucleus, follow it in its orbit and dispatch a lander to the surface. The first spectrospcopic detection of organics on the surface of a comet was made by the Rosetta VIRTIS near-IR spectrometer instrument. Carboxylic acid was suggested as a primary organic material on the nucleus (Capaccioni et al., 2015; Quirico et al., 2016). The Rosetta COSIMA instrument is a secondary ion mass spectrometer equipped with a dust collector, a primary ion gun, and an optical microscope for determining the position and characteristics of the dust particles (Le Roy
et al., 2012). Many organic compounds were identified (Le Roy et al., 2015), and approximately 45% of the refractories in comet 67P/C-G could be in organics (Bardyn
et al., 2017).
@@
Comet C/2013 US (Catalina), a dynamically new Oort cloud comet with a close-Earth approach, was determined to have a 47% amorphous carbon mass fraction (Woodward et al., 2021).
Ultra-carbonaceous Micrometeorites (UCAMMs) represent a small fraction of interplanetary dust particles (IDPs) reaching the Earth, and are exceedingly carbon rich with respect to other types of IDPs (Thomas et al., 1993, 1994), as well as meteoritic samples (Dartois
et al., 2018). They are recovered in the antarctic continent in a relatively pristine state (Duprat
et al., 2010). The bulk fraction of organic refractory material in UCAMMs ranges between 22-58% (Dobrică
et al., 2012). A recent review (Wooden
et al., 2017) pieces together a line of arguments that advocate in favor of a cometary origin for UCAMMs: (a) their crystalline and amorphous silicates are similar to those in comet dust (Wooden, 2008); (b) high S and low Si abundances in glass with embedded metal and sulfides promote a cometary origin (Bradley and
Ishii, 2008); (c) their fluffy textures appear minimally altered by heating during relatively low-velocity atmospheric entry usually associated with cometary origin (Duprat
et al., 2010); (d) they have D/H ratio enrichment expected in cometary dust (Duprat
et al., 2010), including that analyzed by Rosetta (Paquette
et al., 2021); (e) they are globally unequilibrated, in a similar manner as other cometary materials (e.g from Stardust samples (Brownlee et al., 2007)). All this taken together indicates a cometary-like nature for UCAMMs. We also note briefly that many similar arguments also place the origin of relatively carbon-rich chondritic anhydrous IDPs in comets (Engrand
et al., 2023).
@@ 3 Expected composition of Uranus and Neptune
We have previously developed (Podolak et al., 2022) and here further modified a statistical computer code that creates random, monotonic density distributions, which fit a given mass, radius, and moment of inertia. For Uranus and Neptune the gravitational moments J and J have also been measured (Jacobson, 2009, 2014). However, in order to compute these moments for a given density profile, the rotation period of the body must be known, and for both Uranus and Neptune there is some ambiguity in this matter (Helled et al., 2011). In addition, the computation of these moments is extremely time-consuming, so we have chosen, instead, to fit the moment of inertia (MOI) only. While less accurate than using J and J, the difference in the derived density distribution is not significant for our purpose (Podolak et al., 2022).
@@
By integrating the equation of hydrostatic equilibrium we can find the pressure for each density. Podolak et al. (2022) describe several algorithms for associating a temperature and composition with each pressure-density point. Here we use a modified version of the ’inward’ algorithm. Briefly, we assume that at any point the material is composed of a mixture of two adjacent components from the following set of six materials, ordered according to intrinsic density: Fe, SiO, CO, HO, CH, and a solar mix hydrogen/helium gas (hereafter called envelope). Thus SiO can mix with either Fe or CO (or HO for models which do not include carbon compounds), but not with anything else. In addition, the mass fraction of the denser component cannot decrease going inwards. Podolak et al. (2022) did not consider CO and CH, and we now allow for these species as well due to the importance of carbon in outer solar system composition. Starting with a given surface temperature and pressure, we compute the mass fractions of envelope and CH (or HO for models which do not include carbon compounds). Continuing inwards, we try to minimize the mass fraction of the denser material while keeping the temperature constant. If there is no mix that matches under those conditions, we increase the temperature while keeping the same composition. The ’inward’ algorithm tries to maximize the rock to water ratio while minimizing the central temperature. Nevertheless, the central temperatures that are generated can sometimes be quite high, and we discard those models where it exceeds K, which we consider to be unlikely based on models of heat transport in these planets (Podolak et al., 2019; Vazan and Helled, 2020).
@@
For computing random models with Fe (iron), SiO (silica) and HO (water) we use the quotidian equation of state (More et al., 1988; Vazan et al., 2013). For H+He (envelope) material we use SCvH equation of state (Saumon et al., 1995) as modified in (Chabrier et al., 2019). For CO (carbon-monoxide) we use the quotidian equation of state Podolak et al. (2023). For CH (methane) we use the well-known SESAME equation of state Lyon and Johnson (1992). For mixing of different materials we use the additive volume law.
@@ 3.1 Water-rich Models
Newly generated results in this study confirm previous results (Podolak et al., 2022) (and are therefore not shown here explicitly), indicating that the rock/water mass ratio of both Uranus and Neptune is severely restricted when typical materials are considered, which include Fe, SiO, HO, and H+He envelope gas. The exact upper limit value depends sensitively on which moment of inertia factor was assigned to Uranus and Neptune. Whereas classical models considered a MOI factor of 0.23 for both planets (de Pater and Lissauer, 2015), the most recent studies (Neuenschwander and Helled, 2022) suggest that 0.22 is probably a more appropriate value for Uranus and 0.24 for Neptune. We now considered these newer values as lower and upper limits, respectively. Using these new values, the rock/water mass ratio obtained statistically is likewise never more than 2, and actually rarely exceeds 1. In contrast, the water/rock mass ratio has no upper bound, and there are a multitude of water-rich models with far lower central temperatures. These results are consistent with past studies, which almost never found large rock/ice mass ratios.
@@ 3.2 Rock-rich Models
It should be noted that, perhaps, more rock-rich interiors are possible using future algorithms which will consider other material mixing options. Whereas classic layered models usually did not consider mixing at all, we have considered mixing of multiple adjacent materials, while some studies (Helled et al., 2011; Vazan and Helled, 2020) also considered mixing of SiO and H+He, or triple mixing of SiO, HO and H+He (Vazan et al., 2020). Despite these various options, neither Vazan and Helled (2020) nor any other study has ever reached a rock/water mass ratio larger than 2, with one exception. (Helled et al., 2011) managed to accomplish that by suggesting a mixture of pure SiO with a very large fraction (15%) of H+He. While a rock-rich interior is therefore technically possible, recent discussion regarding the question of miscibility (Nettelmann et al., 2016; Miguel and Vazan, 2023) supports the adjacent mixing of SiO and HO, while mixing of SiO and hydrogen (And also Fe and hydrogen, see Wahl et al. (2013); Komabayashi (2021)) remains empirically under-explored and probably requires high temperatures. While our conservative choice of adjacent mixing only is a general limitation, other modes of mixing are perhaps more uncertain and require fine-tuned conditions to obtain a ’desirable’ outcome that leads to rock-rich interiors.
@@ 3.3 Models with Carbon Monoxide
Assuming that CO was indeed prevalent when Uranus and Neptune formed, and also ignoring chemistry and additional concerns, which are addressed later in the paper, we explore the space of possible compositions for Uranus and Neptune given two ices - water and carbon monoxide. If the CO can replace the HO in these models, it should raise the rock/water mass ratio above 2. For comparison purposes, we use an identical numerical setup as in our previous study of randomly generated models (Podolak et al., 2022), only now instead of 4 materials (iron, silica, water and H+He gas), we investigate 5 materials, with carbon monoxide as an additional ice. We consider both Uranus and Neptune and account for MOI factors in the range 0.22-0.24, as previously outlined.
Using the inward algorithm to match 100000 randomly generated density profiles, we show the interior models of Uranus and Neptune with CO in Figure 1. Each random model is depicted by a single pixel, whose color corresponds to the central temperature. The mass fraction of envelope in the planet is shown as a function of the water/rock mass ratio. We note that by rock we refer to the high-Z components SiO+Fe, and by ice we refer to HO+CO (distinguishing ice/rock from water/rock). Since here the water/rock mass ratio extends to very large values, we obtain a better view when the x-axis is in logarithmic scale and truncated at a water-dominated mass ratio of 10.
@@
The dashed vertical line denotes equal proportions of water and rock. The right (left) dotted vertical line indicates a water/rock mass ratio of 2 (0.5). Models with zero rock ( on the x-axis) are not shown. As can be seen, much of the model parameter space is in the region of large water/rock mass ratios. As was the case for the models without CO, there are no models with a water/rock mass ratio less than 0.5, that is, the rock/water mass ratio is never more than 2, and rarely exceeds 1. The reason for this is that, in addition to replacing some HO, the CO predominantly replaces SiO, as can be seen in Figure 2.
@@
CO has a density between that of water and SiO. Therefore the compositional layers transition smoothly from water to CO and then SiO. If the same density profile is fit first without CO ice and then including it, CO can replace either water or SiO. In the models CO tends to replace more of the SiO, than the water. Thus, in Figure 2, we see the mass fraction of each component as a function of radius for a particular Uranus model, originating from the top left of Figure 1, panel (a). Here the MOI is 0.22, the central temperature reaches 12000 K and the water/rock mass ratio is 0.7. Comparing the model with CO (right panel) to that without (left panel) we see that the CO tends to replace the SiO in preference to the water. Thus, simply adding CO to models is not a good solution for the composition tension problem of Uranus and Neptune.

### v2

@@ 1 Introduction
The first models of Uranus and Neptune were constructed based on the premise that they were formed in the outer parts of a primitive proto-planetary nebula, where the elemental abundances were considered solar. Thus, assuming chemical equilibrium, early studies used calculations of the expected relative molecular abundances to gauge the potential composition of the planets that formed in this region (Reynolds and Summers, 1965; Podolak and Cameron, 1974). Ices, such as H2O, CH4, and NH3 were collectively found to be about twice as abundant as rocky materials. Hereafter, we refer to rocks as those refractory minerals in which the ratio of C to silicate-forming elements such as Si, Fe, Mg or Ca is less than 0.1, whereas in refractory organics (CHONs) this ratio is larger than 10 (Fomenkova et al., 1992). Refractory organics are a subset of organic compounds, however which sublimate at higher temperatures, typically in excess of 400 K (Nazari et al., 2023).
These early models typically considered a layered structure, with a rock core, an ice mantle, and an envelope rich in hydrogen and helium (Podolak and Reynolds, 1984; Podolak et al., 1995). Rocks were modeled mainly as silica (SiO2), but sometimes mixed with other molecules (e.g., MgO, Fe, Ni) and water was considered the most abundant and therefore primary ice. The gas envelope was modeled mainly as a H+He solar mix with an admixture of H2O, CH4, and NH3. Later developments highlighted the role of a more complex internal structure including material mixing and gradients (Helled et al., 2020; Helled and Fortney, 2020).
There is increasing evidence however that the constituent building blocks of the outer planets might be refractory-rich, rather than water-rich, challenging common convention with respect to Uranus and Neptune. We summarize these arguments below, indicating that refractory materials in the outer solar system could be 3-6 times more abundant than water.
@@
Large KBOs like Pluto are of such considerable size, that they contain very little residual porosity. Thus, its measured bulk density of 1854 kg m-3 (Stern et al., 2015) directly translates into a similarly rock-rich composition. Other large KBOs like Eris and Haumea often have even higher densities, which imply yet richer rocky compositions (although these are typically associated with further mantle-striping modifications via giant collisions (Lissauer and Stewart, 1993; Leinhardt et al., 2010; Barr and Schwamb, 2016)).
Comets are nowadays regarded as highly refractory-rich bodies. There is a growing body of evidence that strongly suggests comets have refractory to ice mass ratios of at least 3, and potentially even twice that value (Rotundi
et al., 2015; Fulle et al., 2016, 2017, 2019; Choukroun et al., 2020).
Polluted white dwarfs are astrophysical mass spectrometers. Their atmospheres probe the composition of accreted exo-Solar planetary debris, inferring the bulk chemical composition of the pollutants (Hollands et al., 2018; Harrison et al., 2018, 2021). The typical finding is of almost purely rocky debris (Jura and Xu, 2012; Jura and Young, 2014; Pietro Gentile Fusillo et al., 2017). It must be born in mind however that planetesimals can originate from different regions of the disk, devolatilize before their host star becomes a white dwarf (Malamud and Perets, 2016, 2017a, 2017b), and the progression of accretion can complicate the association between pollution and pollutant (Brouwers et al., 2022). Despite these complications, relatively rare few individual cases point to polluting debris which probably originate in the outer parts of their systems and which contain a significant mass fraction of water (Farihi et al., 2013; Raddi et al., 2015; Xu et al., 2017). Yet even in these outlier cases the water fraction is 26%-38%, showing that a large rock/water mass ratio might be a ubiquitous feature of outer planetary disks.
@@ 2 Expected composition of outer solar system planetesimals
Present-day comets typically contain tiny amounts of CO and other hypervolatiles, relatively little water, and are primarily composed of refractory materials (Bockelée-Morvan and Biver, 2017; Fulle et al., 2017, 2019; Choukroun et al., 2020). Observations of various proto-planetary discs indirectly imply that when sufficiently far out in the disc, carbon is depleted from the gas phase and instead enters into solid grains locked inside planetesimals (McClure, 2019; Tabone et al., 2023). Below we discuss direct observations of cometary dust from our own solar system, pointing towards a composition rich in carbon-bearing organic refractories. We show that between 1/3-1/2 of the mass of their refractory material is contained in organic rather than silicate material:
@@
Comet P/Halley was explored by Giotto, a flyby mission in 1986 (Keller et al., 1986), showing that cometary dust is extremely rich in organics relative to typical chondrites, and more similar to dust in dense molecular clouds. High-velocity collisions of dust particles with metal targets on the spacecraft provided a source of ions for mass-spectrometric analysis (Kissel et al., 1986; Lawler and Brownlee, 1992). Particles containing higher fractions of organics were relatively more abundant closer to the nucleus (Fomenkova, 1999), since organic components are more easily dissociated in the coma (Fomenkova et al., 1994; Wooden et al., 2007), providing an extended source for various gaseous species (Cottin and Fray, 2008). Studies also discuss the detailed distribution of carbon among different compounds (Fomenkova, 1999). Overall, the mass ratio of rocks to organics in dust is between 2 and 1 (Fomenkova et al., 1992; Fomenkova, 1999); i.e, carbon-rich organics constitute as 33-50% of the dust mass.
@@
Comet 67P/Churyumov-Gerasimenko was explored by ESA’s Rosetta mission, which was the first to successfully rendezvous with a comet nucleus, follow it in its orbit and dispatch a lander to the surface. The first spectrospcopic detection of organics on the surface of a comet was made by the Rosetta VIRTIS near-IR spectrometer instrument. Carboxylic acid was suggested as a primary organic material on the nucleus (Capaccioni et al., 2015; Quirico et al., 2016). The Rosetta COSIMA instrument is a secondary ion mass spectrometer equipped with a dust collector, a primary ion gun, and an optical microscope for determining the position and characteristics of the dust particles (Le Roy et al., 2012). Many organic compounds were identified (Le Roy et al., 2015), and approximately 45% of the refractories in comet 67P/C-G could be in organics (Bardyn et al., 2017).
@@
Comet C/2013 US10 (Catalina), a dynamically new Oort cloud comet with a close-Earth approach, was determined to have a 47% amorphous carbon mass fraction (Woodward et al., 2021).
Ultra-carbonaceous Micrometeorites (UCAMMs) represent a small fraction of interplanetary dust particles (IDPs) reaching the Earth, and are exceedingly carbon rich with respect to other types of IDPs (Thomas et al., 1993, 1994), as well as meteoritic samples (Dartois et al., 2018). They are recovered in the antarctic continent in a relatively pristine state (Duprat et al., 2010). The bulk fraction of organic refractory material in UCAMMs ranges between 22-58% (Dobrică et al., 2012). A recent review (Wooden et al., 2017) pieces together a line of arguments that advocate in favor of a cometary origin for UCAMMs: (a) their crystalline and amorphous silicates are similar to those in comet dust (Wooden, 2008); (b) high S and low Si abundances in glass with embedded metal and sulfides promote a cometary origin (Bradley and Ishii, 2008); (c) their fluffy textures appear minimally altered by heating during relatively low-velocity atmospheric entry usually associated with cometary origin (Duprat et al., 2010); (d) they have D/H ratio enrichment expected in cometary dust (Duprat et al., 2010), including that analyzed by Rosetta (Paquette et al., 2021); (e) they are globally unequilibrated, in a similar manner as other cometary materials (e.g from Stardust samples (Brownlee et al., 2007)). All this taken together indicates a cometary-like nature for UCAMMs. We also note briefly that many similar arguments also place the origin of relatively carbon-rich chondritic anhydrous IDPs in comets (Engrand et al., 2023).
@@ 3 Expected composition of Uranus and Neptune
We have previously developed (Podolak et al., 2022) and here further modified a statistical computer code that creates random, monotonic density distributions, which fit a given mass, radius, and moment of inertia. For Uranus and Neptune the gravitational moments J2 and J4 have also been measured (Jacobson, 2009, 2014). However, in order to compute these moments for a given density profile, the rotation period of the body must be known, and for both Uranus and Neptune there is some ambiguity in this matter (Helled et al., 2011). In addition, the computation of these moments is extremely time-consuming, so we have chosen, instead, to fit the moment of inertia (MOI) only. While less accurate than using J2 and J4, the difference in the derived density distribution is not significant for our purpose (Podolak et al., 2022).
@@
By integrating the equation of hydrostatic equilibrium we can find the pressure for each density. Podolak et al. (2022) describe several algorithms for associating a temperature and composition with each pressure-density point. Here we use a modified version of the ’inward’ algorithm. Briefly, we assume that at any point the material is composed of a mixture of two adjacent components from the following set of six materials, ordered according to intrinsic density: Fe, SiO2, CO, H2O, CH4, and a solar mix hydrogen/helium gas (hereafter called envelope). Thus SiO2 can mix with either Fe or CO (or H2O for models which do not include carbon compounds), but not with anything else. In addition, the mass fraction of the denser component cannot decrease going inwards. Podolak et al. (2022) did not consider CO and CH4, and we now allow for these species as well due to the importance of carbon in outer solar system composition. Starting with a given surface temperature and pressure, we compute the mass fractions of envelope and CH4 (or H2O for models which do not include carbon compounds). Continuing inwards, we try to minimize the mass fraction of the denser material while keeping the temperature constant. If there is no mix that matches under those conditions, we increase the temperature while keeping the same composition. The ’inward’ algorithm tries to maximize the rock to water ratio while minimizing the central temperature. Nevertheless, the central temperatures that are generated can sometimes be quite high, and we discard those models where it exceeds K, which we consider to be unlikely based on models of heat transport in these planets (Podolak et al., 2019; Vazan and Helled, 2020).
@@
For computing random models with Fe (iron), SiO2 (silica) and H2O (water) we use the quotidian equation of state (More et al., 1988; Vazan et al., 2013). For H+He (envelope) material we use SCvH equation of state (Saumon et al., 1995) as modified in Chabrier et al. (2019). For CO (carbon-monoxide) we use the quotidian equation of state (Podolak et al., 2023). For CH4 (methane) we use the well-known SESAME equation of state (Lyon and Johnson, 1992). For mixing of different materials we use the additive volume law.
@@ 3.1 Water-rich Models
Newly generated results in this study confirm previous results (Podolak et al., 2022) (and are therefore not shown here explicitly), indicating that the rock/water mass ratio of both Uranus and Neptune is severely restricted when typical materials are considered, which include Fe, SiO2, H2O, and H+He envelope gas. The exact upper limit value depends sensitively on which moment of inertia factor was assigned to Uranus and Neptune. Whereas classical models considered a MOI factor of 0.23 for both planets (de Pater and Lissauer, 2015), the most recent studies (Neuenschwander and Helled, 2022) suggest that 0.22 is probably a more appropriate value for Uranus and 0.24 for Neptune. We now considered these newer values as lower and upper limits, respectively. Using these new values, the rock/water mass ratio obtained statistically is likewise never more than 2, and actually rarely exceeds 1. In contrast, the water/rock mass ratio has no upper bound, and there are a multitude of water-rich models with far lower central temperatures. These results are consistent with past studies, which almost never found large rock/ice mass ratios.
@@ 3.2 Rock-rich Models
It should be noted that, perhaps, more rock-rich interiors are possible using future algorithms which will consider other material mixing options. Whereas classic layered models often did not consider mixing at all (especially in the rock component), we have considered mixing of multiple adjacent materials, while some studies (Helled et al., 2011; Vazan and Helled, 2020) also considered mixing of SiO2 and H+He, or triple mixing of SiO2, H2O and H+He (Vazan et al., 2020). Despite these various options, neither Vazan and Helled (2020) nor any other study has ever reached a rock/water mass ratio larger than 2, with one exception. Helled et al. (2011) managed to accomplish that by suggesting a mixture of pure SiO2 with a very large fraction (15%) of H+He. While a rock-rich interior is therefore technically possible, recent discussion regarding the question of miscibility (Nettelmann et al., 2016; Miguel and Vazan, 2023) supports the adjacent mixing of SiO2 and H2O, while mixing of SiO2 and hydrogen (And also Fe and hydrogen, see Wahl et al. (2013); Komabayashi (2021)) remains empirically under-explored and probably requires high temperatures. While our conservative choice of adjacent mixing only is a general limitation, other modes of mixing are perhaps more uncertain and require fine-tuned conditions to obtain a ’desirable’ outcome that leads to rock-rich interiors.
@@ 3.3 Models with Carbon Monoxide
Assuming that CO was indeed prevalent when Uranus and Neptune formed, and also ignoring chemistry and additional concerns, which are addressed later in the paper, we explore the space of possible compositions for Uranus and Neptune given two ices - water and carbon monoxide. If the CO can replace the H2O in these models, it should raise the rock/water mass ratio above 2. For comparison purposes, we use an identical numerical setup as in our previous study of randomly generated models (Podolak et al., 2022), only now instead of 4 materials (iron, silica, water and H+He gas), we investigate 5 materials, with carbon monoxide as an additional ice. We consider both Uranus and Neptune and account for MOI factors in the range 0.22-0.24, as previously outlined.
Using the inward algorithm to match 100000 randomly generated density profiles, we show the interior models of Uranus and Neptune with CO in Figure 1. Each random model is depicted by a single pixel, whose color corresponds to the central temperature. The mass fraction of envelope in the planet is shown as a function of the water/rock mass ratio. We note that by rock we refer to the high-Z components SiO2+Fe, and by ice we refer to H2O+CO (distinguishing ice/rock from water/rock). Since here the water/rock mass ratio extends to very large values, we obtain a better view when the x-axis is in logarithmic scale and truncated at a water-dominated mass ratio of 10.
@@
The dashed vertical line denotes equal proportions of water and rock. The right (left) dotted vertical line indicates a water/rock mass ratio of 2 (0.5). Models with zero rock ( on the x-axis) are not shown. As can be seen, much of the model parameter space is in the region of large water/rock mass ratios. As was the case for the models without CO, there are no models with a water/rock mass ratio less than 0.5, that is, the rock/water mass ratio is never more than 2, and rarely exceeds 1. The reason for this is that, in addition to replacing some H2O, the CO predominantly replaces SiO2, as can be seen in Figure 2.
@@
CO has a density between that of water and SiO2. Therefore the compositional layers transition smoothly from water to CO and then SiO2. If the same density profile is fit first without CO ice and then including it, CO can replace either water or SiO2. In the models CO tends to replace more of the SiO2, than the water. Thus, in Figure 2, we see the mass fraction of each component as a function of radius for a particular Uranus model, originating from the top left of Figure 1, panel (a). Here the MOI is 0.22, the central temperature reaches 12000 K and the water/rock mass ratio is 0.7. Comparing the model with CO (right panel) to that without (left panel) we see that the CO tends to replace the SiO2 in preference to the water. Thus, simply adding CO to models is not a good solution for the composition tension problem of Uranus and Neptune.